Suppose a young couple deposits $600 at the end of each quarter in an account that earns 7.2%, compounded quarterly, for a period of 4 years. How much is in the account after the 4 years? (Round your answer to the nearest cent.)

After the 4 years, they start a family and find they can contribute only $200 per quarter. If they leave the money from the first 4 years in the account and continue to contribute $200 at the end of each quarter for the next 18 1 2 years, how much will they have in the account (to help with their child's college expenses)? (Round your answer to the nearest cent.) $
Answer #1

Answer page; 01 initially p= $ 600 per quarter 1.7.20% per quarter => 710/4 = 0.072 = 0.018 :-0.018 na 4x4 = 16 quarters So fpage:02 600 (1.33034-13 10.0721 = 600 (0.33034) 10.072) = 600 ( 0.33034) 10.018) - Goox 18.35222 = Ifol. 33 s -$11,0ll. 33 afpage :03 = 41226,394 = 41226. 39 P = 000 per quositer :- 7.24.14 per quarter no 18.5x4 = 14 quenter S2 = future value for fwi